OpenStudy (anonymous):

10 divided by 2/5 = ?

OpenStudy (anonymous):

4 12 15 25

OpenStudy (alexandervonhumboldt2):

\[10/\frac{ 2 }{ 5 }=10*\frac{ 5 }{ 2 }=5*5=25\]